OpenStudy (anonymous):

Need help explain mew for integrating factors on 1st order linear derivatives. for example dv/dt = 9.8 - 0.196v

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Integrating_factor

OpenStudy (bahrom7893):

basically you try to find an integrating factor, such that, when both sides are multiplied by it, the left side turns into a derivative of something, which makes it easy to integrate.

OpenStudy (anonymous):

rewrite your equation as: v' = 9.8 - 0.196v v' + 0.196v = 9.8 or in a more general sense, v' + p(x)v = q(x), then \[\rho = e^ \left( \int\limits p(x) dx \right)\] or mew, whichever symbol you're using. Than if you multiply your integrating factor through the equation, you get a simple derivation on the left, namely (mew*v)' = mew*q(x) Then you integrate and win